### v2

# Zhiyuan Fan     Yumeng Wang     Sandeep Polisetty     Yi R. (May) Fung

Abstract
Large Vision Language Models (LVLMs) excel in various vision-language tasks. Yet, their robustness to visual variations in position, scale, orientation, and context that objects in natural scenes inevitably exhibit due to changes in viewpoint and environment remains largely underexplored. To bridge this gap, we introduce V2R-Bench, a comprehensive benchmark framework for evaluating Visual Variation Robustness of LVLMs, which encompasses automated evaluation dataset generation and principled metrics for thorough robustness assessment. Through extensive evaluation on 21 LVLMs, we reveal a surprising vulnerability to visual variations, in which even advanced models that excel at complex vision-language tasks significantly underperform on simple tasks such as object recognition. Interestingly, these models exhibit a distinct visual position bias that contradicts theories of effective receptive fields, and demonstrate a human-like visual acuity threshold. To identify the source of these vulnerabilities, we present a systematic framework for component-level analysis, featuring a novel visualization approach for aligned visual features. Results show that these vulnerabilities stem from error accumulation in the pipeline architecture and inadequate multimodal alignment. Complementary experiments with synthetic data further demonstrate that these limitations are fundamentally architectural deficiencies, scoring the need for architectural innovations in future LVLM designs. [^1] Our code and data will be available at https://github.com/toward-agi/Visual-Variations-Robustness.
[Uncaptioned image]  V2R-Bench: Holistically Evaluating LVLM Robustness to Fundamental Visual Variations
Zhiyuan Fanthanks: Equal contribution.     Yumeng Wang     Sandeep Polisetty     Yi R. (May) Fung Hong Kong University of Science and Technology zhiyuan.fan@connect.ust.hk
Figure 1: Illustration of V2R-Bench. Left: Automated data generation framework that adds visual variations to images, covering basic visual tasks, existing benchmarks, as well as supplementary visual tasks. Right: Metrics for robustness testing, including performance consistency, semantic and token-level stability, and LLM-as-a-judge.
1 Introduction
The rapid development of Large Vision Language Models (LVLMs) (Liu et al., 2023; Lu et al., 2024a) has been driven by two key factors: innovations in model architectures and the availability of high-quality training data. These models have demonstrated impressive results in complex vision-language tasks, achieving human-level performance across various challenges (Fei et al., 2024). To systematically evaluate such capabilities, numerous multimodal benchmarks have been developed to assess models’ fundamental knowledge (Fu et al., 2024b), perceptual capability (Wu and Xie, 2023; Zhang et al., 2024b), cognitive understanding (Fu et al., 2024a), and reasoning skills (Lu et al., 2024b; Huang et al., 2024) across various downstream applications.
While current benchmarks extensively evaluate models using images collected in specific scenarios, they overlook a more fundamental generalization capability: the robustness of LVLMs to visual variations. In input images, objects naturally exhibit diverse variations: spatial positions shift with changes in camera angles and viewpoints; object scales vary depending on viewing distances; orientations deviate from standard poses through rotations and inversions; and objects appear in a range of visual semantic contexts. These visual variations raise concerns about LVLMs’ robustness: whether models maintain consistent perception capabilities across all spatial positions in input images; what visual acuity threshold determines reliable performance; and how changes in orientation and visual context influence model behavior.
Despite the importance of such robustness, current research has primarily focused on model robustness to socio-cultural factors (Ananthram et al., 2024), adversarial prompt attacks (Wu et al., 2024b; Liu et al., 2024b), or corrupted image inputs (Liu et al., 2024a), leaving the impact of natural visual variations largely unexplored.
In this paper, we propose V2R-Bench, a comprehensive benchmark framework to evaluate LVLM robustness against fundamental visual variations. Our framework consists of an automated data generation pipeline, and tailored evaluation metrics, which can readily extend to various VQA tasks. Through extensive evaluation on 21 LVLMs, we uncover surprising findings: despite their excellence in complex multimodal tasks, these models exhibit unexpected vulnerabilities to visual variations, leading to poor performance even in basic tasks such as object recognition. Specifically, we observe: (1) a counter-intuitive position bias where models achieve higher accuracy at image edges rather than the center; (2) a human-like visual acuity threshold where model reliability steadily decreases with object size, reaching and maintaining minimum performance below a critical scale threshold; (3) selective robustness to certain orientations while remaining fragile to others; and (4) a tendency to ground predictions on visual contextual inference rather than direct visual perception.
To identify the root cause of these vulnerabilities, we conduct a systematic analysis of LVLM components. A key innovation in our analysis is a novel visualization method that reconstructs language tokens from aligned visual features, offering insights into how models process and transmit visual information across modalities. Our investigation reveals that inadequate multimodal alignment is the primary bottleneck, as models fail to maintain stable visual representations across variations and struggle to effectively align visual semantics with the language model. To disentangle whether these limitations stem from architectural constraints or data deficiency, we conduct complementary experiments with synthetic training data. The results reveal that these vulnerabilities are fundamentally rooted in architectural design. These findings underscore two critical directions for future LVLM development: stronger multimodal alignment mechanisms to maintain semantic consistency across variations, and unified architectural designs to mitigate error accumulation in current pipeline structures.
Our contributions are summarized as follows:
  • We identify and formulate a novel problem: the robustness of LVLMs to visual variations, a fundamental yet overlooked capability essential for reliable vision-language reasoning.
  • We propose V2R-Bench, an evaluation framework with automated data generation and tailored metrics, uncovering significant vulnerabilities in current LVLMs through an extensive evaluation of 21 models.
  • We develop a systematic component-level analysis with a novel visualization technique for multimodal alignment, revealing the root causes of these vulnerabilities and providing insights for future architectural improvements.
2 Related Work
Large Vision Language Models.
The remarkable progress of LVLMs stems from continuous advances in model architectures and the exponential growth of training datasets. On the architectural front, the rapid progress in foundation language models (OpenAI et al., 2024; Touvron et al., 2023; Jiang et al., 2023; Bai et al., 2023a) and vision models (Radford et al., 2021; Oquab et al., 2024), together with increasingly sophisticated multimodal alignment modules (Liu et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Zhu et al., 2023; Wang et al., 2024b; Zhou et al., 2024), has established the fundamental capabilities of LVLMs in understanding and reasoning across modalities. Building upon these architectural foundations, the emergence of high-quality datasets (Chen et al., 2023a; Zhao et al., 2023; Wang et al., 2023b; Li et al., 2023c) designed for different training stages (e.g., multimodal alignment, visual instruction tuning, preference alignment) has enabled LVLMs to demonstrate exceptional performance across diverse real-world scenarios. More recently, research efforts have focused on optimizing lightweight model architectures and curating datasets tailored to edge scenarios (Ma et al., 2024; Chu et al., 2024; Zhu et al., 2024; Yao et al., 2024; Lee et al., 2025), thereby promoting the practical deployment and adoption of LVLMs in real-world applications.
While the pipeline architecture of these LVLMs effectively leverages pretrained domain-specific knowledge from vision and language components, it potentially accumulates errors and vulnerabilities across modules. However, no systematic investigation has been conducted to attribute model failures to individual architectural components, limiting both model interpretability and understanding of modality alignment.
LVLM Evaluation and Benchmarking.
Recent years have witnessed the emergence of numerous benchmarks for evaluating LVLM capabilities across cognitive and perceptual dimensions (Liang et al., 2024; Fu et al., 2024b, a; Liu et al., 2024d; Yu et al., 2024; Li et al., 2023a; Lee et al., 2024; Zhang et al., 2025), assessing various aspects including reasoning skills, understanding abilities, and inherent knowledge. The scope of evaluation has further extended into specialized domains, with an increased emphasis on real-world scenario performance, as researchers develop dedicated benchmarks for embodied intelligence (Wang et al., 2024a; Yang et al., 2023; Zhang et al., 2023), medical image analysis (Xia et al., 2024; Chen et al., 2024a; Hu et al., 2024), chart understanding Huang et al. (2025), and robotic control (Chen et al., 2024b; Wang et al., 2023c). In parallel with capability assessment, researchers have begun investigating LVLM robustness from two critical perspectives: first, examining semantic biases in model responses, particularly those related to societal factors like gender and racial prejudices (Wang et al., 2024c; Howard et al., 2024; Steed and Caliskan, 2021; Wu et al., 2024a); and second, analyzing adversarial vulnerabilities via carefully crafted visual prompts for assessing model reliability under targeted attacks (Liu et al., 2024a; Luo et al., 2024b; Zhang et al., 2024a; Wang et al., 2024c; Chen et al., 2023b). Yet, the capacity to withstand basic visual variations, a fundamental aspect of robustness widely present in real-world deployment, remains unexplored in current research.
3 Evaluation Framework
In this section, we present our V2R-Bench evaluation framework, which includes an automated pipeline for generating diverse visual variations across task settings (3.1), our evaluation protocols (3.2), and a detailed description of the construction and characteristics of our evaluation dataset (3.3).
3.1 Automated Variations Generation
The automated data generation pipeline incorporates four fundamental visual variations that are prevalent in real-world scenarios. Position variations investigate whether LVLMs exhibit blind spots in their visual processing, where models may fail to accurately perceive visual information at certain spatial locations. Meanwhile, scale variations examine the perceptual boundaries of LVLMs when processing objects at different scales, similar to clinical vision tests for humans. Orientation variations challenge the ability of LVLMs to process objects at different rotational angles, which is crucial for real-world scenarios like robotic navigation where objects rarely appear in canonical orientations. Moreover, context variations test whether model predictions remain consistent across diverse environmental settings, to reveal whether LVLMs perform genuine visual perception or rely primarily on contextual cues for inference.
Formally, given an image $I$, a set of transformed images along these dimensions is generated as:
$\mathcal{D}=\{T(I,v)|v\in\{P\times S\times R\times C\}\}$ (1)
where $P=\{1,...,W\}\times\{1,...,H\}$, $S=[s_{min},s_{max}]$, $R=[0,2\pi]$, and $C\in\mathcal{B}$ represent the sets of position, scale, rotation, and context variations respectively. For each original question-image pair, this generation process produces $|P|\times|S|\times|R|\times|C|$ variants in total, enabling a holistic exploration of the entire variation space that leaves no potential vulnerability unexplored.
3.2 Evaluation Metrics
In assessing LVLM robustness, we consider performance consistency, which measures whether a model maintains its task-specific metrics across visual variations, by quantifying how these variations impact model performance:
$C_{m}(I)=1-\sqrt{\frac{1}{N}\sum_{v\in\mathcal{V}}(M(I_{v})-\overline{M})^{2}}$ (2)
where $M(I_{v})$ denotes the task-specific metric on variation $I_{v}$, $\overline{M}$ represents the mean performance across all variations, and $\mathcal{V}$ is the set of variations. A larger $C_{m}$ indicates better consistency, with 1 representing that the model is robust enough to be unaffected by visual variations.
We also evaluate the output stability of model generation, at both semantic and token levels:
$S_{s}(I)=\frac{1}{|\mathcal{V}|^{2}}\sum_{v_{i}\in\mathcal{V}}\sum_{v_{j}\in\mathcal{V}}sim(E(O_{v_{i}}),E(O_{v_{j}}))$ (3)
$S_{t}(I)=\frac{1}{|\mathcal{V}|^{2}}\sum_{v_{i}\in\mathcal{V}}\sum_{v_{j}\in\mathcal{V}}\frac{|T(O_{v_{i}})\cap T(O_{v_{j}})|}{|T(O_{v_{i}})\cup T(O_{v_{j}})|}$ (4)
where $E(O)$ denotes the output embedding, $sim(\cdot,\cdot)$ computes the cosine similarity between embeddings, and $T(O)$ represents the set of tokens in output $O$.
Furthermore, LLM-as-a-judge (Zheng et al., 2023) is employed to emulate human assessment of LVLM-generated outputs under structured visual variations, providing an additional qualitative perspective on model robustness.
3.3 Dataset Construction
The proposed automated generation pipeline is implemented across two categories of tasks. The first category examines fundamental visual capabilities through object recognition and direction recognition tasks, with target objects and directional indicators being systematically transformed through our visual variations using image processing algorithms and inpainting diffusion models (Corneanu et al., 2024; Lugmayr et al., 2022)
The second category extends to existing multimodal benchmarks, focusing on scenarios where variations in position, scale, orientation, and context preserve ground-truth validity, ensuring that any performance changes reflect model robustness rather than ground-truth alteration. The final evaluation datasets contain a total of 428K images. Each category serves a distinct yet complementary purpose in our evaluation: the basic tasks provide controlled, interpretable measures of fundamental capabilities, while the extended benchmarks assess robustness in more naturalistic settings. The detailed implementation of these generation algorithms is provided in Appendix C.
Model (Accuracy) (Robustness)
Position Orientation Scale Context Position Orientation Scale Context
Obj. Dir. Obj. Dir. Obj. Dir. Obj. Dir. Obj. Dir. Obj. Dir. Obj. Dir. Obj. Dir.
Qwen-VL 3.7 28.5 2.3 28.1 15.3 47.7 5.1 47.7 91.3 89.7 92.4 77.3 89.1 89.9 83.2 87.9
Qwen2-VL-7B 5.9 40.2 6.7 42.7 17.5 50.9 6.7 42.7 94.5 91.0 95.5 66.4 86.2 92.3 93.7 89.0
Molmo-7B-D 26.0 62.6 26.2 64.6 27.8 63.2 26.2 64.6 91.1 91.0 95.5 66.4 86.2 92.3 93.7 89.0
H2OVL-2B 5.5 39.8 6.3 42.3 17.2 50.6 6.3 42.3 83.3 91.0 97.7 100 92.6 100 100 97.9
H2OVL-800M 1.3 20.1 1.1 19.9 17.9 33.2 17.3 30.8 79.3 89.0 92.1 74.5 83.2 92.6 89.5 90.1
Phi3-Vision 1.5 36.7 1.8 37.9 6.4 37.5 1.8 37.9 96.8 92.9 100 62.2 93.7 100 100 92.9
Phi3.5-Vision 6.2 40.8 7.0 41.7 7.9 43.0 6.7 41.7 95.5 92.3 94.5 63.5 100 100 93.7 92.3
InternVL-Mono 5.7 40.0 6.5 42.5 17.7 51.0 6.5 42.5 91.5 90.5 93.5 74.5 91.5 95.5 92.5 91.5
InternVL-2 6.1 40.5 6.9 43.0 17.9 51.3 6.9 43.0 94.5 92.5 96.5 78.5 94.5 97.5 95.5 94.5
InternVL-2.5 5.8 40.3 6.6 42.8 17.6 51.1 6.6 41.8 96.5 93.5 98.5 81.5 96.5 98.5 97.5 96.5
LLaVA1.5 5.6 39.9 6.4 42.4 17.4 50.8 6.4 42.4 95.5 92.5 97.5 76.5 95.5 98.5 96.5 95.5
LLaVA-Onevision-Qwen2 16.0 44.6 17.3 45.9 24.5 47.4 17.3 45.9 92.9 91.1 89.5 71.2 92.9 100 83.0 90.5
LLaVA1.6-mistral 1.7 19.0 1.9 19.4 4.4 21.6 1.9 19.4 96.8 93.7 100 78.8 96.8 96.8 100 96.8
LLaVA1.6-vicuna 1.8 19.2 2.0 19.6 4.5 21.8 2.0 19.6 96.5 93.5 98.5 77.5 96.5 97.5 98.5 96.5
LLaMA3.2-11B-V 5.7 40.0 6.5 42.5 17.7 51.0 6.5 42.5 92.5 91.3 94.5 75.2 92.5 96.5 93.0 92.5
LLaMA3.2-90B-V 6.3 41.2 7.1 43.7 18.2 51.7 7.1 43.7 97.5 94.5 98.6 82.5 97.5 98.4 98.5 97.5
GLM4-V-9B 14.1 27.9 9.3 11.7 38.1 25.7 1.2 23.3 95.3 92.5 97.1 79.6 95.9 97.0 96.1 95.5
MiniCPM-V-2 5.7 39.2 1.3 39.3 15.6 41.9 6.3 42.8 93.2 91.2 95.2 76.2 93.7 96.3 94.8 93.1
MiniCPM-V-2.5 9.2 44.8 9.0 48.3 21.0 55.5 9.0 46.3 95.3 92.2 97.3 78.2 95.7 97.2 96.8 94.8
MiniCPM-V-2.6 5.9 40.5 6.7 43.0 17.8 51.3 6.7 43.0 96.2 93.3 98.2 79.3 96.8 98.3 97.7 95.9
GPT-4o 31.5 78.2 29.8 77.2 31.9 69.5 26.7 79.9 98.3 96.2 100 85.3 98.7 100 100 98.2
Table 1: Comprehensive evaluation of model performance (i.e., accuracy and robustness) under different visual variations, for object and direction detection tasks with metrics defined in Section 3.2.
4 Cross-Modality Diagnosis
To understand the underlying mechanisms of LVLM vulnerabilities to visual variations, we propose a systematic analysis framework that examines contributions of each model components to robustness issues. Central to this framework is a novel visualization technique that provides straightforward insights into how visual features extracted by vision encoder are processed through the multimodal alignment module and aligned with language embedding space.
4.1 Component-level Analysis
Here we first formalize the general architecture and process pipeline of modern LVLMs to facilitate subsequent analysis. Given an input question $Q$ and an image $I$, the vision encoder $E_{v}$ first extract visual features from the input image:
$v=E_{v}(I;\theta_{v})\in\mathbb{R}^{N_{v}\times D_{v}}\vspace{-0.4em}$ (5)
These visual features are then projected into the language embedding space through a multimodal alignment module $P$:
$h=P(v;\theta_{p})\in\mathbb{R}^{N_{v}\times D_{h}}\vspace{-0.4em}$ (6)
Finally, the language model $M$ takes the aligned visual features in conjunction with question embeddings $E_{l}(Q)\in\mathbb{R}^{N_{q}\times D_{h}}$ as input to generate the response in an autoregressive manner:
$R_{t}=M(h,E_{l}(Q),R_{<t};\theta_{m})\quad\text{for }t=1,\dots,T$ (7)
As the initial input module of LVLMs, the vision encoder $E_{v}$ fundamentally determines the performance ceiling of the whole model, since the visual information contained in its extracted features represents the upper bound of visual content available to subsequent modules. Given the distinct pretraining paradigms of vision encoders (contrastive learning (Radford et al., 2021) and self-supervised learning (Oquab et al., 2024)), the feature quality assessment differs accordingly: for contrastive-trained encoders, the corresponding text encoder enables zero-shot analysis, while for supervised-trained ones, linear probing (Alain and Bengio, 2018) and clustering analysis serve to assess their extracted features.
The multimodal projector $P$ acts as a bridge between the visual feature and language embedding spaces, which raises two questions for its performance analysis: (RQ1) Do the projected features $h$ preserve the visual information contained in $v$?, and (RQ2) Do the projected features $h$ align well with the embedding space of the language model $M$? To answer these questions, two analytical approaches are employed: (1) comparing the performance of pre-projection features $v$ and post-projection features $h$ on visual tasks to quantify potential visual information loss during multimodal alignment, and (2) measuring the discriminability between projected visual features $h$ and the language embeddings of their corresponding captions to assess the quality of modality alignment.
Figure 2: (a) Demonstration of position bias effects.  (b) Accuracy heatmaps for object recognition and direction recognition, across object scales and position variations.  (c) Model accuracy as a function of relative object scale.
The language model $M$, as the final module for integrating aligned visual features with input questions and generating responses, is inherently affected by potential errors from upstream modules, which complicates the assessment of intrinsic language model robustness. Thus, an evaluation strategy is devised to bypass upstream modules by directly providing visual information as language tokens, which simulates ideal visual feature extraction and multimodal alignment where visual information is perfectly preserved without any loss or misalignment, enabling a controlled setting for analyzing inherent language model capabilities. We create text-based evaluation datasets mirroring LVLMs’ visual tasks, using matrix-structured text to simulate ideally encoded visual scenes with diverse variations. Comparing performance on text-based versus visual tasks identifies if language models cause vulnerabilities to visual variations.
4.2 Visual-Linguistic Feature Analysis
In addition to quantitative analysis of multimodal alignment, our proposed framework also incorporates a novel visualization approach, which reconstructs language tokens from aligned visual features to provide interpretable evidence of the alignment process, an aspect previously unexplored. Specifically, given a single aligned visual feature $h\in\mathbb{R}^{1\times D_{h}}$ and token embedding matrix $E\in\mathbb{R}^{|V|\times D_{h}}$ of the language model $M$, the feature $h$ can be decoded to a set of language tokens that approximate its semantic meaning:
$t=\text{topk}\left(\text{softmax}\left(hE^{\top}\right)\right),\vspace{-0.55em}$ (8)
where $|V|$ denotes the language model vocabulary size and $k$ controls the number of selected tokens.
Through these decoded language tokens, the semantic meaning captured by the aligned visual features becomes intuitively understandable, which helps address the following questions: (RQ3) how do language models interpret these aligned visual features? (RQ4) How robust is the semantics of aligned visual features to visual variations? These questions are investigated in detail in Section  5.2.
5 Experiments
Detailed experiment settings are provided in Appendix D.
5.1 Impact of Visual Variations
The evaluation results across different LVLM archi-
tectures and model scales are presented in Tables 1. Despite their impressive performance demonstrated on complex visual tasks, these models exhibit surprising vulnerability to simple visual variations, resulting in significantly degraded performance across basic visual tasks. Even proprietary models such as gpt-4o, Claude, and Gemini not only produce incorrect outputs, but also occasionally claim an inability to perceive the visual content entirely, as illustrated in Appendix F. Interestingly, despite claims that some carefully distilled smaller models outperform their larger counterparts on mainstream benchmarks, our analysis reveals that scaling laws still hold for robustness: within the same model architecture, larger models consistently demonstrate better stability across visual variations.
For an in-depth understanding of these vulnerabilities, this subsection focuses on LLaVA model, which provides full access to its training data, code, and parameters. Position: Figure 2(b) shows that model performance varies dramatically across different positions. Contrary to the effective receptive field theory (Luo et al., 2017; Raghu et al., 2022) which suggests vision models have better perception of central regions, LVLMs exhibit strong visual position bias with higher accuracy at peripheral regions (as shown in Table 2), raising concerns about their fundamental visual processing mechanisms. Figure 2(a) presents an intuitive example using right-pointing arrows as input, illustrating the significant effect of positional variations on prediction outcomes. Scale: Figure 2(c) illustrates a sharp performance decline as object size decreases,
Model Object Direction
Middle Surrounding Middle Surrounding
LLaVA1.6 1.94 1.95 (+0.01) 17.94 19.99 (+2.05)
LLaVA1.5 1.83 1.89 (+0.06) 15.33 18.02 (+2.69)
Vila 1.87 1.93 (+0.06) 16.85 17.83 (+0.98)
Ovis 2.07 2.11 (+0.04) 23.77 31.08 (+7.31)
InternVL 2.01 2.79 (+0.78) 21.58 30.26 (+8.68)
Qwen2-VL 5.78 7.02 (+1.24) 39.36 44.08 (+4.72)
LLaMA3.2-11B-V 3.72 6.58 (+2.86) 27.19 35.88 (+8.69)
Table 2: Quantitative analysis of position bias. Experimental results show that these models consistently demonstrate better perception of surrounding regions compared to the central areas.
stabilizing when the object occupies 1/100 of the image area (equivalent to 1/10 of both width and height). This reveals a potential visual acuity threshold in LVLMs, similar to human vision, that defines a critical boundary below which model outputs become unreliable, serving as a key indicator for deploying LVLMs in fine-grained visual perception tasks. Orientation: performance degradation is observed across different orientations, with models exhibiting distinct directional biases: some orientations show robustness, while others lead to significant failures. Interestingly, models like Fuyu and BLIP demonstrate a pronounced predictive tendency, being heavily influenced by the left orientation. Context: model predictions vary with different contextual arrangements, correlating with context content, which raises questions about whether the model truly perceives the target objects or infers them from contextual cues.
To further investigate the root causes of these vulnerabilities, we design complementary tasks focusing on fundamental perceptual capabilities. To examine whether position and direction vulnerabilities stem from limitations in spatial and directional perception, we introduce two specialized tasks: coordinate identification, where models directly output position coordinates to test spatial understanding, and path tracking, where models sequentially output coordinates following directional lines to examine continuous directional perception. To verify whether models achieve genuine recognition or merely rely on contextual inference, we develop a modified OCR task where certain letters in fluent text are intentionally replaced with incorrect characters and then blurred at different levels, testing whether models faithfully report the visual content.
As shown in Table 3, all models exhibit poor performance in coordinate recognition and path tracing, indicating these vulnerabilities fundamentally originate from the inability to accurately perceive spatial properties. Figure 3 reveals peak accuracy at initial points followed by monotonically decreasing performance, with marginal recovery at terminal nodes. This indicates not only weak isolated direction recognition but also inadequate performance in visual-following reasoning tasks, suggesting that spatial reasoning capabilities likely derive more from commonsense world knowledge rather than genuine visual understanding. Table 6 results further demonstrate that LVLMs tend to overconfidently output contextually inferred conclusions rather than objectively depicted visual content, while LLMs achieve better accuracy under the same task conditions. This contrast proves that the observed limitations stem not from denoising characteristics (Devlin et al., 2019) inherent to the transformer architecture itself, but rather from interference introduced by visual modality integration.
Model EMA PM-IA PM-SA PA
Qwen2-VL 6.0 23.9 7.9 48.7
Molmo-7B-D 0.0 16.0 5.3 19.9
Phi3-Vison 6.2 44.8 18.8 70.7
Phi3.5-Vison 0.5 9.3 3.0 12.3
LLaVA-Onevision-Qwen2 0.5 7.4 2.1 7.7
LLaVA1.6 0.0 1.7 0.5 1.3
Table 3: Path and coordinate task accuracy in %, of 6 selected models across metrics as defined in Appendix C.2.
Figure 3: Performance evaluation on a 6-point path tracing task, where accuracy indicates coordinate prediction precision at each sequential point.
5.2 Component Analysis
Vision Encoder.
Figure 4 reveals that changes in the prediction label probabilities of the vision encoder align with shifts in the next token logits of the LVLM when scale and context variations are introduced. This behavioral consistency suggests that LVLM inherits vulnerability to visual variations from its vision encoder component.
Figure 4: Changes in vision encoder classification probabilities and LVLM token predictions under different visual variations (context and scale). For each variation, we show the ViT’s top-3 class probabilities (left) and LVLM’s top-5 token logits (right), demonstrating how semantic interpretations shift across visual variations.
Multimodal Projector.
The following analyses address the four research questions posed above.
(RQ1) Visual Information Loss. Linear probing results for pre-projection and post-projection features are presented in Table 4. The significant performance degradation after multimodal projection (i.e., the MM-Projector) suggests irrecoverable semantic information loss during the modality alignment process, contributing to LVLM’s vulnerability across visual variations. We further compare the fine-tuned vision encoder in LLaVA 1.6 with the original one. While its vision encoder is fine-tuned to adapt image-text features for generative tasks, its linear probing performance underperforms the original vision encoder, revealing a trade-off: task-specific adaptation improves multimodal coherence but erodes the vision encoder’s innate spatial representational fidelity. This aligns with LLaVA 1.5’s limitations, where only the lightweight projector can be fine-tuned and struggles to mitigate information loss from the frozen CLIP encoder. Our results underscore that while recent advancements in alignment neural networks alleviate alignment bottlenecks, fundamental architectural constraints such as patch-based tokenization and positional bias persist as critical vulnerabilities, necessitating unified approaches to relax the conflicts between visual grounding and multimodal alignment.
Model ViT-LVLM MMProj-LVLM
Obj. Dir. Obj. Dir.
LLaVA1.5 44.2 82.6 3.2   (-41.0) 11.7 (-70.9)
LLaVA1.6 40.6 91.2 14.1 (-26.5) 20.7 (-70.5)
Vila 21.9 96.5 9.7   (-12.2) 13.2 (-83.3)
InternVL 22.1 96.4 13.2 (-8.9) 33.9 (-62.5)
Ovis 21.9 96.5 13.3 (-8.6) 31.8 (-64.7)
Qwen2-VL 44.6 100 65.9 (+21.3) 100  (0)
LLaMA3.2-11B-V 22.2 100 11.9 (-10.3) 98.6 (-1.4)
Table 4: Linear probing accuracy (%) of vision language models on Object (Obj.) and Direction (Dir.) tasks reveals the collaborative impact of vision encoders/multimodal projectors and LVLMs. Notably, Qwen2-VL, which employs a native vision encoder trained directly with the language model rather than pretraining on image modality first and then aligning to language modality, shows no performance loss and even improvement after modality projection, suggesting the potential of unified architectures.
Figure 5: Comparison of different feature representation visualizations. (a) Visualization of visual features, aligned features and text embeddings. (b) Word cloud of decoded features. (c) Visual-linguistic token alignment.
(RQ2) Inadequate Multimodal Alignment. Figure 5(a) illustrates the spatial distribution of visual features, aligned features, and language embeddings. We observe that the distribution of the aligned features is similar to that of image features, meaning that the multimodal process preserves the visual semantics for some extent. Ideally, the aligned features should closely resemble the language embeddings. However, our observation reveals a significant disparity between the aligned feature space and the language embedding space, suggesting a lack of adequate modality alignment. This finding underscores the critical role of the multimodal projector, which is a key factor contributing to the LVLM’s vulnerability to visual variations. To explore the image feature representation with regard to visual variations, Figure 8 in the appendix presents the clustering analysis of image features of the same object at different directions and positions. These variations introduce substantial alterations in image features, highlighting the inherent challenges faced by robust image encoding techniques. These fluctuations can lead to inconsistencies in feature extraction and representation across different instances of the same object or scene. The limited robustness in current encoding methodologies may struggle to effectively capture and encode these variations while maintaining the semantic integrity of the features, potentially compromising the accuracy and reliability of downstream tasks. Addressing this issue requires the development of more resilient encoding strategies that can adapt to diverse visual transformations, enhancing the overall robustness and generalizability of image feature representations in complex visual tasks.
(RQ3) Interpretation of Aligned Features. Figure 5(b) demonstrates the results of decoding aligned visual features into language tokens. Due to the inherent differences in attention patterns between vision encoder (bidirectional) and language model (autoregressive), these decoded results do not form coherent natural language, with only a subset of tokens being semantically relevant to the image content. These aligned visual features reside outside the discrete language embedding space, echoing back to the visualization shown in Figure 5(a). This can be interpreted as visually conditioned soft prompts (Lester et al., 2021; Gu et al., 2022; Liu et al., 2022), providing implicit cues that contain image-related information to guide model responses to image-related queries, see Appendix E. However, prior work (Bailey et al., 2024) has revealed the limitations of such soft prompt approaches, including their instability and bugs due to lack of interpretability and the discrepancy from language model embedding space, ultimately leading to the vulnerability to visual variations.
(RQ4) Visual Semantics Vulnerability. Based on the results for RQ3, significant changes in decoded language tokens are observed across visual variations, indicating poor semantic robustness, as shown in Figure 5(c). These inconsistent soft prompt prefixes consequently lead to substantial fluctuations in model outputs, further compromising model robustness.
Language Model.
It has come to the community’s attention that positional biases in natural language widely exist in language models Wang et al. (2023a); Jung et al. (2019); Zheng et al. (2023); Koo et al. (2024), where models tend to favor information appearing earlier in the text. As suggested in Table 5, we observe that language models exhibit position bias in matrix-structured text representations of visual information. Although object and background information are explicitly represented as language tokens, model predictions are still influenced by context, albeit to a lesser extent than LVLMs, suggesting that the vulnerability primarily stems from the upstream vision encoder and multimodal alignment module, where the aligned visual features fed into the language model are already heavily influenced by visual context. The low accuracy in the coordinate task demonstrates that, similar to LVLMs, language models also lack precise positional awareness of objects. These limitations are rooted in the autoregressive nature of language models, which tends to prioritize sequential dependencies over structural relationships.
Model Number Coordinate Object
w/o BG w/ BG w/o BG w/ BG w/o BG w/ BG
LLaMA3-8B 6.88 10.12 0.32 0.24 99.96 84.64
LLaMA3.1-8B 17.42 13.06 0.80 0.27 97.54 91.89
Mistral-v0.2 29.40 7.76 0.80 0.00 88.32 85.28
Qwen-2 32.88 13.64 0.88 0.84 99.52 97.84
Qwen-2.5 10.19 8.11 5.45 1.99 99.44 97.19
Deepseek-7B 18.26 16.33 1.06 0.93 85.43 41.71
Table 5: Performance of LVLM language model backbones on text-based tasks: comparing scenarios with tokenized background represented as asterisks (w/ BG) versus random words (w/o BG).
6 Mitigating Robustness Issues
To determine whether the robustness issues stem from architectural limitations or insufficient training data, we explore improvements through two complementary approaches: First, we conduct controlled experiments on a subset of our test data while maintaining a held-out test set, directly probing the architectural capacity for robust visual understanding. Second, we utilize a general visual instruction tuning dataset injected with visual variations to analyze whether a more diverse training distribution can enhance model robustness.
We find that a more diverse dataset offers limited improvement in model robustness, likely due to insufficient data volume (as exhaustively covering these variations requires many more variants). Moreover, while directly training the model on spatial visual tasks improves its performance on position and direction tasks, it does not enhance robustness against position and orientation variations. As Figure 6 suggests, after training, despite improvements in the coordinate task, the path tracing task remains underperforming, due to the lack of sustained visual attention. This reveals that the vulnerability is fundamentally rooted in architectural design choices rather than data limitations.
Figure 6: The comparison of benchmark evaluation results with upper bound on coordinate and path tasks, where the upper bound is the testing result of the LVLM fine-tuned on partial benchmark dataset.
The design of the vision encoder utilizes patch-based tokenization (Radford et al., 2021) and positional embeddings, which may lead to information fragmentation due to arbitrary patch partitioning and position sensitivity induced by explicit positional encoding. Furthermore, the cascading pipeline architecture amplifies vulnerabilities at each component, underscoring the need for a new architectural approach. Although unified architectures currently underperform on visual understanding tasks due to training stability challenges (Chen et al., 2025a), this direction remains promising.
7 Conclusion
We present V2R-Bench, an evaluation framework designed to assess the robustness of LVLMs against visual variations. Our results show significant vulnerabilities in existing LVLMs and identify their origins as twofold: insufficient multimodal alignment and error accumulation inherent in pipeline model architectures. While synthetic data augmentation Qin et al. (2025) seems as a mitigation strategy, fundamental advancement in the field requires a shift towards native multimodal architectures rather than the current approach of concatenating separate language and vision modalities. We aim to draw greater attention to the importance of visual robustness in LVLMs and inspire future research toward more robust architectural paradigms.
Limitations
Aligned with the paper track’s focus, this work primarily concentrates on identifying a novel problem, establishing corresponding evaluation methods, and providing initial analytical insights into these vulnerabilities. Regarding solutions to the identified vulnerabilities, only two data synthesis approaches were explored, while potential improvements through architectural modifications, pre-training strategies, test-time self-correction or self-consistency approaches He et al. (2024); Wang et al. (2025), and novel modality alignment methods remain unexplored due to their substantial resource requirements. The detailed analysis presented in this paper aims to provide insights for future research addressing these challenges. Additionally, while this study examines four fundamental types of visual variations, it does not exhaust the infinite possibilities of visual transformations, leaving some long-tail cases unexplored.
Ethics Statements
This study reveals a fundamental vulnerability in LVLMs - their lack of robustness to visual variations that naturally occur from camera parameter adjustments and environmental changes. Such robustness deficiency results in significant output inconsistencies across visually similar scenarios, compromising model reliability in real-world deployments.
A critical security implication emerges from these findings: visual variations could serve as a novel attack vector. Through strategic object placement with specific positions, scales, orientations, or contexts, these fundamental visual variations could be exploited to manipulate model behavior. Unlike conventional adversarial attacks requiring sophisticated training procedures, this approach requires no training and generates natural images without artificial artifacts, making such attacks particularly challenging for existing detection mechanisms.
The identified vulnerability underscores the need for increased attention within the research community to these fundamental yet profoundly impactful visual variations, rather than solely pursuing state-of-the-art performance on complex tasks. Enhanced robustness to these variations is crucial for ensuring consistent model performance in natural environments. Moreover, as attacks based on visual variations exploit inherent model vulnerabilities rather than crafted adversarial strategies, addressing this robustness issue becomes essential for improving both the reliability and security of deployed LVLMs.
Appendix A Discussion
Summary of Empirical Findings
Through the evaluation of V2R-Bench, we discover that introducing visual variation causes inconsistent and unstable output. Such vulnerability is due to the error accumulation in the data generation pipeline. Figure 7 reveals that ViT has the tendency to depend on the context during inference rather than truly recognize a target object, which is similar to LVLM. Through component analysis, we claim that multimodal projector is the main cause of the vulnerability to visual variations. The inadequate multimodal alignment causes visual information loss and the decoding result of aligned feature does not form coherent natural language.
The Use of Synthetic Data
While our framework enables the incorporation of real-world benchmarks, we acknowledge that generated variations may not fully capture the complexity and diversity of real-world visual inputs. To address this, future work can be done to further integrate our framework with diffusion-based contextual blending or 3D synthetic data generation technique such as Blender-SDG Arenas (2020).
Future Directions
V2R-Bench fills a crucial gap in existing evaluations by systematically testing robustness to fundamental visual variations (position, scale, orientation, context)—ubiquitous in real-world scenarios but overlooked by current benchmarks. For instance, autonomous vehicles require consistent object recognition regardless of camera angles, and medical imaging tools must identify anomalies across scales. By integrating existing benchmarks and evaluating with our synthetic data, our framework directly addresses these needs by enabling models to be stress-tested under realistic conditions. Researchers can also easily expand tests to new variations (e.g., lighting changes) or domains (e.g., CT image).
The benchmark’s component-level analysis offers fine-grained diagnostic information, guiding researchers toward targeted improvements. A key future direction lies in enhancing the multimodal alignment mechanisms. Additionally, the identified vulnerabilities are not merely the result of limited data or training strategies but also stem from fundamental architectural constraints requiring further investigation.
Appendix B Prompt Template
Prompt For Evaluation Object: Identify the object in the image. Direction: List the direction the arrow is pointing in the image using one of the following: up, down, left, right, top-left, bottom-left, top-right, or bottom-right. Coordinate: This is a coordinate plot with a single point. Provide the coordinate in the format $(x,)$ for 1D, $(x,y)$ for 2D, or $(x,y,z)$ for 3D. Path: Describe the coordinates of each point along the line from the start to the end in the format $[(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})]$.
Model Visual Dependency Knowledge Dependency
Text B0 B1 B2 B3 Text B0 B1 B2 B3
GPT-4o 0.2109 0.0475 0.0586 0.2916 0.4724 0.0821 0.2044 0.2030 0.2004 0.2232
Qwen2-VL 0.2531 0.1169 0.2417 0.3710 0.5344 0.2674 0.2344 0.2284 0.3154 0.4945
LLaVA-1.6 0.2129 0.5732 1.8404 1.8026 1.4573 0.3743 0.5676 1.8013 1.7919 1.4485
Table 6: Performance Comparison of Models under Different Blur Conditions: LVLMs favor contextual inference over input-output consistency, while LLMs maintain better alignment with inputs. As blur levels increase, LVLMs confidently continue contextual reasoning, implying they inherently operate through inference from ambiguous visual signals rather than direct visual processing.
Figure 7: Linear probing results for Vision Encoder showing position bias patterns similar to those observed in LVLMs.
Figure 8: t-SNE visualization of aligned features under directional and positional variations, demonstrating partial sensitivity to visual variations in the feature space.
Appendix C Experimental Setup
We conduct comprehensive experiments across three distinct datasets to evaluate our model’s performance under various conditions. The experimental settings for each dataset are detailed below.
C.1 Coordinate Dataset
For the coordinate dataset, we systematically vary four key parameters to thoroughly assess model performance. The point range parameter defines the spatial extent of the coordinate system, with values spanning from confined spaces ($[-5,5]$) to broader ranges ($[-10,10]$, $[0,10]$, and $[0,20]$). This variation allows us to evaluate the model’s ability to handle different scales of spatial information.
To investigate the impact of visual aids on model performance, we experiment with both the presence and absence of reference lines and grid systems. These binary parameters (True/False) help us understand how additional visual context affects the model’s coordinate understanding.
Regarding dimensionality, we focus our analysis on one- and two-dimensional coordinate systems. While three-dimensional coordinates were initially considered, they were ultimately excluded from our final experiments due to consistently poor model performance across preliminary tests.
C.2 Path Dataset
In the path dataset experiments, we explore the model’s capability to process and understand connected point sequences. We vary the complexity of paths by adjusting the number of points from 2 to 6, creating a progression from simple linear paths to more complex multi-point trajectories. The spatial distribution of these points is controlled through the same range parameters as the coordinate dataset: $[-5,5]$, $[-10,10]$, $[0,10]$, and $[0,20]$. For each unique combination of point count and range setting, we generate a substantial set of 100 images, ensuring robust evaluation across different configurations.
We evaluate the performance of path tracing and coordinate recognition through the following metric:
(1) Exact Match Accuracy (EMA)
A predicted path is considered correct only if it exactly matches the ground truth answer.
(2) Partial Match Order-Independent Accuracy (PM-IA)
In this metric, a point in the predicted path is deemed correct if its coordinates match any point in the ground-truth path. PM-IA is computed as the average accuracy across all positions.
(3) Partial Match Order-Sensitive Accuracy (PM-SA)
This metric marks a point in the predicted path as correct only if both its coordinates and position in the path match those of a point in the ground truth path. PM-SA is calculated as the average accuracy across all positions.
(4) Point Accuracy (PA)
This metric is the accuracy of the coordinate recognition.
C.3 Object and Orientation Dataset
The object and orientation dataset is designed to evaluate model understanding of object positioning and directionality. We carefully select ten distinct object categories, including eight animals (shiba dog, cat, bear, eagle, snake, panda, turtle, and fish) and two vehicles (car and plane). This diverse selection allows us to assess model performance across varying object morphologies and complexities.
Scale perception is tested through a comprehensive set of object-to-background ratios: 1/2, 1/3, 1/5, 1/10, 1/15, and 1/20. These ratios represent a wide spectrum from prominent objects (1/2) to more subtle presentations (1/20). Furthermore, we evaluate each object against two distinct background types: solid colors for controlled conditions and semantic images for real-world complexity.
The orientation aspect of our experiments encompasses eight distinct directions: the four cardinal directions (up, down, left, right) and their intermediates (top-left, bottom-left, top-right, bottom-right). This comprehensive directional coverage allows us to assess the model’s ability to understand and interpret various object orientations.
Through these carefully designed experimental settings, we aim to provide a thorough and systematic evaluation of our model’s capabilities across different aspects of visual understanding and spatial reasoning.
C.4 Text Dataset
We create char matrices of size 8*8, 16*16, 24*24, 32*32, 40*40 and 64*64. A target word is selected from one of the following, [‘dog’, ‘cat’, ‘bird’, ‘lion’, ‘tiger’, ‘zebra’, ‘monkey’, ‘panda’], and is positioned within the matrices. Except the object, the rest of the matrices are either asterisks (e.g. corresponding to the w/o BG setting) or random background words. We design 3 tasks to test the robustness of LLM in the cross-component analysis: (1) target word recognition, corresponding to object detection task in image; (2) coordinate recognition which corresponds to coordinate recognition; and (3) object counting, which is a fundamental skill needed in path tracing. By systematically varying the background contexts while maintaining the target object, we assess whether models truly recognize objects independently or merely rely on contextual associations.
Appendix D Experimental Setup
D.1 Evaluated Models
Qwen-VL
The Qwen-VL model family represents Alibaba’s cutting-edge vision-language model series. The family includes three main variants: the original Qwen-VL Bai et al. (2023b), which established the foundation for vision-language processing; Qwen2-VL-7B, offering a balanced mid-size option; and Qwen2-VL-72B, the largest and most sophisticated version featuring state-of-the-art visual understanding capabilities and support for videos over 20 minutes long Wang et al. (2024b). The latest addition, QVQ-72B-Preview, serves as an experimental research model specifically focused on advancing visual reasoning capabilities Team (2024).
Molmo
Molmo-7B-D Deitke et al. (2024) is an open-source vision-language model developed by the Allen Institute for AI, built on Qwen2-7B and utilizing OpenAI’s CLIP as its vision backbone.
H2OVL
The H2OVL-Mississippi Galib et al. (2024) model family are specifically designed for efficient on-device applications and privacy-focused use cases. The family consists of two specialized models: H2OVL-Mississippi-0.8B, a compact model optimized for text recognition that achieves state-of-the-art performance on OCRBench, and H2OVL-Mississippi-2B, a model for general vision-language tasks including image captioning and visual question answering.
Phi-3
Microsoft’s Phi-3-vision and Phi-3.5-vision Abdin et al. (2024) represent a significant advancement in multimodal AI. Phi-3.5-vision, the latest iteration, is a lightweight yet powerful model featuring a 128K token context length and support for both single and multi-image processing.
InternVL
The InternVL family includes: Mono-InternVL Luo et al. (2024a), which established the foundation with its vision-language capabilities; InternVL-2 Chen et al. (2024c), which expanded the model sizes and improved performance; and InternVL-2.5 Chen et al. (2025b), the latest iteration that introduces significant architectural and training improvements.
LLaVA
LLaVA-1.5 Liu et al. (2023) uses pretrained CLIP Radford et al. (2021) and Vicuna language model as the backbone, establishing the foundation with impressive performance across 12 benchmark datasets. LLaVA-OneVision-Qwen2 Li et al. (2024) pushes performance boundaries across single-image, multi-image, and video scenarios while enabling strong task transfer capabilities. LLaVA-1.6Liu et al. (2024c), or LLaVA-NeXT, further enhances capabilities with increased input resolution, improved visual reasoning, and enhanced OCR capabilities.
LLaMA
LLaMA 3.2 Vision models  Dubey et al. (2024) represent Meta’s latest advancement in multimodal AI, introducing vision capabilities to the LLaMA family for the first time. The 11B and 90B parameter versions are specifically designed to handle both text and image inputs, featuring a novel architecture that integrates image encoder representations into the language model.
GLM-4
GLM-4V-9B GLM et al. (2024) supports high-resolution image processing at 1120*1120 pixels and enables dialogue capabilities in both Chinese and English.
MiniCPM-V
MiniCPM-V Yao et al. (2024) is a series of multimodal large language models (MLLMs) designed specifically for deployment on end-side devices like mobile phones and personal computers.
D.2 Implementation Details
Our experiments are conducted on 8 NVIDIA H100 GPUs. All models maintain their original parameter configurations during inference, with an average processing speed of 525 tokens per second.
To examine spatial reasoning capabilities analogous to those in LVLMs, we design a suite of tasks that evaluate text-based object recognition, counting, and spatial analysis. Using a matrix-structured text input that simulates idealized visual encoding, we assess three fundamental capabilities: (1) target token identification amid background elements, (2) frequency quantification of target tokens, and (3) spatial localization of target tokens within the matrix. The dataset specifications and distributions are detailed in Appendix C.4.
Appendix E Visually Conditioned Soft Prompt
In this section, we explain why aligning visual features to the language model’s embedding space through multimodal alignment module can be viewed as a form of visually conditioned soft prompt. Soft prompting prepends learned vectors to the language model’s input, optimizing these vectors during training to achieve desired tasks. These vector prefixes learn latent instructions during the tuning process:
$\mathbf{h}_{t}$ $=LM([\underbrace{P}_{\text{prompt}};\mathbf{x}]_{<t}),$ (9)
$\text{where }P\in\mathbb{R}^{l\times d},\quad\mathcal{L}=\sum_{t}-\log p(x_{t}|\mathbf{h}_{t})$
where $P$ represents the learned soft prompt vectors and $\mathbf{x}$ is the input token embeddings. In LVLM’s multimodal alignment process, extracted visual features are mapped to the language model’s embedding space through a neural network based multimodal alignment module. The module is trained to transform visual features into embedding space that serve as special task prefixes, guiding model responses to image-related queries::
$\mathbf{h}_{t}$ $=LM([\underbrace{f_{\theta}(E_{v}(I))}_{\text{visual prompt}};\mathbf{x}]_{<t}),$ (10)
where =$\text{where }\mathcal{L}=\sum_{t}-\log p(x_{t}|\mathbf{h}_{t},I)$
where $f_{\theta}$ is the alignment network and $\mathbf{v}$ represents the visual features. Fundamentally, both approaches optimize vector prefixes independent of the discrete language embedding matrix. The key difference lies in their optimization objectives: soft prompts aim to generate task-relevant responses based on the dataset, learning fixed prefixes that remain constant after training. In contrast, while aligned visual features are also trained on the entire dataset, they produce dynamic, image-specific prefixes during inference. Therefore, aligned visual features can be considered a special case of soft prompts - visually conditioned soft prompts - sharing the characteristic of residing in a representation space distinct from the language embedding space.
Appendix F Failure Examples
  • ChatGPT: Sorry, I don’t see the arrow in the image.
  • Claude: A sleek fish glides through the harbor waters, its scales catching glints of sunlight filtering through the surface. The backdrop features wooden pier posts rising from the water and moored fishing boats gently bobbing in the distance. The water has a deep blue-green tint typical of a working harbor, with small ripples creating patterns of light and shadow. The fish appears to be navigating between the shadows cast by the docks, perhaps searching for food or shelter among the port’s underwater structures.
  • Gemini: Reasons:
    • Image Information: The image only shows an animal on the water’s surface, without indicating the specific direction of the animal’s head, body, or tail.
    • Lack of Reference Points: There are no other objects or background elements to help determine the animal’s orientation.
    Therefore, it is impossible to determine the animal’s orientation from the information provided in the image.
    If you can provide more information about the image, such as the shooting angle or the animal’s posture, it may be possible to determine its orientation more accurately.
  • InternVL: This image shows a pure white background without any recognizable objects or text. It is a simple plane with no complex designs or patterns.